# Amorphous Solid Model of Vectorial Hopfield Neural Networks

### v3

@@ Abstract
The Hopfield model Hopfield (1982) is a cornerstone of associative memory in neural networks and artificial intelligence Meshulam and Bialek (2025). In its classical formulation, $N$ binary neurons $s_{i}\in\{-1,+1\}$ evolve according to symmetric interactions $W_{ij}$ defined by the Hebbian rule from stored patterns. The dynamics minimizes $E=-\frac{1}{2}\sum_{ij}W_{ij}s_{i}s_{j}$, leading to pattern retrieval via convergence to energy minima.
@@
We consider 3D networks of $N$ nodes (neurons) connected by 3D vectors (“bonds” or synapses) as in Fig. 1. In the amorphous solid analogy, the connections are springs with spring constant $\kappa$ Milkus and Zaccone (2016, 2017); Zaccone (2023).
Figure 1: Rendering of a 3D network of $N$ nodes with average coordination number $Z=7$ and nearly uniform orientation distribution.
The binary Hopfield model consists of $N$ neurons with $s_{i}\in\{-1,+1\}$. To store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ with $\boldsymbol{\xi}^{\mu}=(\xi_{1}^{\mu},\ldots,\xi_{N}^{\mu})$ and $\xi_{i}^{\mu}\in\{-1,+1\}$, the weight matrix is
@@
In the vectorial model, each neuron is a unit vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ with $|\mathbf{s}_{i}|=1$. We store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$, where each pattern is a set of $N$ random unit vectors on $S^{2}$:
@@
which satisfies $W\Xi=\Xi$, so each stored pattern is an exact eigenvector of $W$ with eigenvalue $1$. This guarantees perfect recall of all training patterns in the fully connected case and provides a natural extension of the classical formulation to correlated patterns.
@@
where $X_{i}$ collects neighbour states and $Y_{i}$ the target states for neuron $i$. The parameter $\lambda$ ensures numerical stability and controls smoothness. This local pseudo-inverse generalizes the fully connected rule to sparse graphs while encoding topology.
@@
The vectorial Hopfield model generates energy landscapes with multiple local minima corresponding to stored patterns. We quantify pattern stability via the energy difference between stored and random configurations. For $N=25$ and $\gamma=0.1$, stored patterns have mean energies about $7$ units lower than random states, ensuring robust retrieval against initialization noise. The energy function
@@
Convergence properties depend on initialization and pattern density. The convergence time scales logarithmically with system size $N$ for successful retrievals, while basin size decreases approximately exponentially as $\gamma$ approaches $\gamma_{c}$, and basin boundaries become more complex. The unit-vector constraint produces dynamics on the product manifold $S^{2}\times S^{2}\times\cdots\times S^{2}$, preserving pattern structure but adding geometric complexity. The update rule

### v4

@@ Abstract
The Hopfield model Hopfield (1982) is a cornerstone of associative memory in neural networks and artificial intelligence Meshulam and Bialek (2025). In its classical formulation, NN binary neurons $s_{i}\in\{-1,+1\}$ evolve according to symmetric interactions $W_{ij}$ defined by the Hebbian rule from stored patterns. The dynamics minimizes $E=-\frac{1}{2}\sum_{ij}W_{ij}s_{i}s_{j}$, leading to pattern retrieval via convergence to energy minima.
@@
We consider 3D networks of NN nodes (neurons) connected by 3D vectors (“bonds” or synapses) as in Fig. 1. In the amorphous solid analogy, the connections are springs with spring constant $\kappa$ Milkus and Zaccone (2016, 2017); Zaccone (2023).
Figure 1: Rendering of a 3D network of NN nodes with average coordination number $Z=7$ and nearly uniform orientation distribution.
The binary Hopfield model consists of NN neurons with $s_{i}\in\{-1,+1\}$. To store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ with $\boldsymbol{\xi}^{\mu}=(\xi_{1}^{\mu},\ldots,\xi_{N}^{\mu})$ and $\xi_{i}^{\mu}\in\{-1,+1\}$, the weight matrix is
@@
In the vectorial model, each neuron is a unit vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ with $|\mathbf{s}_{i}|=1$. We store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$, where each pattern is a set of NN random unit vectors on $S^{2}$:
@@
which satisfies $W\Xi=\Xi$, so each stored pattern is an exact eigenvector of $W$ with eigenvalue 11. This guarantees perfect recall of all training patterns in the fully connected case and provides a natural extension of the classical formulation to correlated patterns.
@@
where $X_{i}$ collects neighbour states and $Y_{i}$ the target states for neuron ii. The parameter $\lambda$ ensures numerical stability and controls smoothness. This local pseudo-inverse generalizes the fully connected rule to sparse graphs while encoding topology.
@@
The vectorial Hopfield model generates energy landscapes with multiple local minima corresponding to stored patterns. We quantify pattern stability via the energy difference between stored and random configurations. For $N=25$ and $\gamma=0.1$, stored patterns have mean energies about 77 units lower than random states, ensuring robust retrieval against initialization noise. The energy function
@@
Convergence properties depend on initialization and pattern density. The convergence time scales logarithmically with system size NN for successful retrievals, while basin size decreases approximately exponentially as $\gamma$ approaches $\gamma_{c}$, and basin boundaries become more complex. The unit-vector constraint produces dynamics on the product manifold $S^{2}\times S^{2}\times\cdots\times S^{2}$, preserving pattern structure but adding geometric complexity. The update rule